OpenStudy (anonymous):

Find the length of the radius of the following circle. (x + 2)2 + y2 = 10

OpenStudy (anonymous):

\[(x-4)^2+(y-k)^2=r^2\] has radius \(r\) in your case \(r^2=10\) and so \(r=\sqrt{10}\)

OpenStudy (anonymous):

thanks